# Optimization under rare events: scaling laws for linear chance-constrained programs

### v1

@@ 7. Proofs
Assume that $c\geq 0$, $g_{\ell}$ is non-decreasing in $y$ for each $\ell$ and that $g_{\ell}\rightarrow g$ pointwise as $\ell\rightarrow\infty$, with $g$ continuous and strictly increasing. Let $v_{*}=\sup\{c^{T}y:g(y)\leq 1,y\geq 0\}$. Further, define $v_{\ell}=\sup\{c^{T}y:g_{\ell}(y)\leq 1,y\geq 0$}. Then, $v_{\ell}\rightarrow v_{*}$.
If $c=0$ the result is immediate. So, we assume that at least one entry of $c$ is positive. We first consider the case $v_{*}=\infty$. For every $a>0$ there exists $y\geq 0$ such that $c^{T}x>a$. By continuity of $g$ there exists $\gamma\in(0,1)$ such that $g(\gamma y)<1$ and $c^{T}y>a/2$. This implies that $g_{\ell}(\gamma y)\leq 1$ and therefore $v_{\ell}\geq a/2$, for all $\ell$ large enough. Thus, we may assume that $v_{*}<\infty$. Next, observe that for each $\eta<1$, $g_{\ell}(\eta y^{*})<1$ for $\ell$ sufficiently large. Consequently, $c^{T}y_{\ell}\geq\eta c^{T}y^{*}$ for $\ell$ sufficiently large, and by letting $\eta\uparrow 1$ we obtain that
$\liminf_{\ell\rightarrow\infty}c^{T}y_{\ell}\geq c^{T}y^{*}.$ (27)
To prove a corresponding asymptotic upper bound, we assume to the contrary that for some $\varepsilon>0$ and a sub-sequence $(v_{\ell})$ (which we also index by $\ell$) we have $v_{\ell}\geq v_{*}+2\varepsilon$.
We want to contradict the optimality of $y^{*}$. Let us partition $c=(\bar{c},0)$, where the part $\bar{c}$ has strictly positive elements. We then note that $v_{*}=\max\{\bar{c}^{T}\bar{y}:g(\bar{y},0)\leq 1,\bar{y}\geq 1\}$. Because $g_{\ell}$ is non-decreasing, we have that $v_{\ell}=\max\{\bar{c}^{T}\bar{y}:g_{\ell}(\bar{y},0)\leq 1,\bar{y}\geq 0\}.$ So, we may assume without loss of generality that $c$ has strictly positive elements. Now, let $y_{\ell}$ be an optimal solution corresponding to $v_{\ell}$ (if the solution is not achieved, we can work with an $\varepsilon/2$ solution). Since $g_{\ell}$ is non-decreasing, we can make the coordinates of $y_{\ell}$ smaller, obtaining $y^{\prime}_{\ell}$ which remains feasible and $c^{T}y^{\prime}_{\ell}=c^{T}y^{*}+\varepsilon$. Since the elements of $c$ can be assumed to be strictly positive, the set $C=\{x:c^{T}y=c^{T}y^{*}+\varepsilon\}$ is compact. Since the $g_{\ell}$’s are non-decreasing, the convergence $g_{\ell}\rightarrow g$ occurs uniformly on compact sets. By choosing $\ell$ sufficiently large we can construct $y^{\prime\prime}_{\ell}$ (possibly by reducing the coordinates of $y^{\prime}_{\ell}$) such that $g(y^{\prime\prime}_{\ell})\leq 1$, contradicting the optimality of $y^{*}$. This completes the proof of this proposition. $\square$
@@ 7.1. Proof of Theorem 1
We first show that the feasible region of ($\rm{LT}$) is convex and compact. For this, it suffices to show that $\{b:I(b)\geq 1\}$ is convex and compact, where $I$ is defined in (9). Let $b^{(1)}$ and $b^{(2)}$ be such that $I(b^{(i)})\geq 1,i=1,2$. Let $a\in(0,1)$. For the desired convexity property, we have to show that $I(ab^{(1)}+(1-a)b^{(2)})\geq 1$. Since $\lambda$, defined in (7), is continuous, non-decreasing, and satisfies the relation $\lambda(rx)=r^{\beta}\lambda(x)$, there exists an $x^{(a)}\geq 0$ such that $I(ab^{(1)}+(1-a)b^{(2)})=\lambda(x^{(a)})$. Since $(ab^{(1)}+(1-a)b^{(2)})x^{(a)}\geq 1$, we have that either $b^{(1)}x^{(a)}\geq 1$ or $b^{(2)}x^{(a)}\geq 1$. Suppose for concreteness that $b^{(1)}x^{(a)}\geq 1$. Then, $I(ab^{(1)}+(1-a)b^{(2)})\geq I(b^{(1)})\geq 1$ implies the desired convexity property. In fact, a similar argument as that provided in the proof of Proposition 1, shows that $I(\cdot)$ is continuous (using that the minimization defining $I(b)$ can be assumed to be taken over a compact set); this implies that $\{b:I(b)\geq 1\}$ is closed. The compactness of the set $\{b:I(b)\geq 1\}$ follows from the relation $\lambda(rx)=r^{\beta}\lambda(x)$, $\beta>0$.
Next, we show that (10) holds, using standard textbook arguments from large-deviation theory, see e.g. Dembo and Zeitouni (1998): note that, if $B_{\delta}(x)$ is the $L_{\infty}$ ball of radius $\delta$ around $x$, the assumption on $L$ implies that $-\log{\bf P}(L/r\in B_{\delta}(x))\sim\inf_{u\in B_{\delta}(x)}\lambda(u)$. Therefore, using also the continuity of $\lambda$, the same estimate holds for the closure of $B_{\delta}(x)$, from which it follows that $L/r$ satisfies a weak large-deviations principle with rate function $\lambda$ and speed $q(r)$. Since $\lambda(ax)=a^{\beta}\lambda(x)$, $\lambda$ has compact level sets. Therefore, $L/r$ satisfies a full large-deviations principle with rate function $\lambda$ and speed $q(r)$. Since $b^{T}L$ is a continuous function of $L$ and $\lambda$ is continuous, (10) follows from the contraction principle.
We now turn to the proof of Theorem 1. We apply Proposition 1 with $\ell=1/\delta$, $g_{\ell}(x)=(\log\delta)/\log{\bf P}(\phi(x,L)>q^{-1}(\log 1/\delta))$, and $g(x)=1/J(x)$. The pointwise convergence $g_{\ell}(x)\rightarrow g(x)$ follows from (10). The fact that $g$ is strictly increasing follows from the homogeneity property of $J$. $\square$
@@ 7.2. Proof of Theorem 2
Since $v_{{\bf CVaR},\delta}\leq v_{\delta}$, it suffices to prove an asymptotic lower bound. For this, we utilize Theorem 1. Let $y^{*}$ be a solution of ($\rm{LT}$) and for fixed $\gamma>0$, define $y:=(1-\gamma)y^{*}$. Moreover, let $r_{\rm{LT},\delta}:=q^{-1}(\log 1/\delta)$. We will prove that $x=y/r_{\rm{LT},\delta}$ will be feasible for ($\rm{CVaR}_{\delta}$), for $\delta$ sufficiently small.
The optimal value $\tau=\tau_{\delta}$ in (5) with $x=y/r_{\rm{LT},\delta}$ solves a first-order optimality condition given by
$\delta={\bf P}\big{(}\phi(y,L)>r_{\rm{LT},\delta}(\tau_{\delta}+1)\big{)}.$ (28)
Since this equality is true for all $\delta>0$, this implies that $r_{\rm{LT},\delta}(\tau_{\delta}+1)\to\infty$. Combining this with the tail behavior of $\phi(y,L)$ given in Equation (11), we see that as $\delta\downarrow 0$
$\log(1/\delta)=(1+o(1))J(y)q\big{(}r_{\rm{LT},\delta}(\tau_{\delta}+1)\big{)},$ (29)
and therefore
$r_{\rm{LT},\delta}(\tau_{\delta}+1)=q^{-1}\Big{(}(1+o(1))\frac{\log(1/\delta)}{J(y)}\Big{)}=J(y)^{-1/\beta}q^{-1}(\log 1/\delta)(1+o(1)).$ (30)
Recalling that $r_{\rm{LT},\delta}=q^{-1}(\log 1/\delta)$ by definition for light-tailed distibutions, we see that
$\tau_{\delta}+1=J(y)^{-1/\beta}(1+o(1)).$ (31)
Now, recall $y=(1-\gamma)y^{*}$, so that $J(y)=(1-\gamma)^{-\beta}J(y^{\ast})>1$ by (11), (9), as $J(y^{\ast})=1$. We conclude that $\tau_{\delta}\rightarrow-\gamma$ as $\delta\downarrow 0$.
Next we use that the constraint in ($\rm{CVaR}_{\delta}$) with $x=y/r_{\rm{LT},\delta}$ is equivalent to
${\bf E}\big{[}\phi(y,L)/r_{\rm{LT},\delta}\mid\phi(y,L)>r_{\rm{LT},\delta}(\tau_{\delta}+1)\big{]}\leq 1.$ (32)
We compute that
${\bf E}[\phi(y,L)/r_{\rm{LT},\delta}$ $\mid\phi(y,L)/r_{\rm{LT},\delta}>\tau_{\delta}+1]$
$\leq 1-\gamma/2+{\bf E}[(\phi(y,L)/r_{\rm{LT},\delta}-(1-\gamma/2))^{+}\mid\phi(y,L)/r_{\rm{LT},\delta}>\tau_{\delta}+1]$
$=1-\gamma/2+\frac{\int_{1-\gamma/2}^{\infty}{\bf P}(\phi(y,L)>r_{\rm{LT},\delta}u)\mathrm{d}u}{{\bf P}(\phi(y,L)>(1+\tau_{\delta})r_{\rm{LT},\delta})}.$
For fixed $y$, we can write ${\bf P}(\phi(y,L)>z)=\exp\{-\tilde{q}(z)\}$, with $\tilde{q}$ regularly varying of index $\beta>0$. Due to the Potter bounds (see e.g. Bingham et al. (1987), Theorem 1.5.6), for every $\varepsilon>0$ there exists a constant $z_{\varepsilon}$ such that, for $u\geq s\geq 1/2$, and $r\geq z_{\varepsilon}$
$\tilde{q}(ru)/\tilde{q}(rs)\geq(1-\varepsilon)(u/s)^{\beta-\varepsilon}.$ (33)
Using this bound, for sufficiently small $\varepsilon$ (w.r.t. $\gamma$ and $\beta$) and recalling that $\tau_{\delta}\rightarrow-\gamma$, it follows that
$\frac{\int_{1-\gamma/2}^{\infty}{\bf P}(\phi(y,L)>r_{\rm{LT},\delta}u)\mathrm{d}u}{{\bf P}(\phi(y,L)>(1+\tau_{\delta})r_{\rm{LT},\delta})}$ $\int_{1-\gamma/2}^{\infty}\exp\big{(}-\tilde{q}(r_{\rm{LT},\delta}u)+\tilde{q}((1+\tau_{\delta})r_{\rm{LT},\delta})\big{)}\mathrm{d}u$
$=\int_{1-\gamma/2}^{\infty}\exp\Big{(}-\tilde{q}((1+\tau_{\delta})r_{\rm{LT},\delta})\Big{(}\frac{\tilde{q}(r_{\rm{LT},\delta}u)}{\tilde{q}((1+\tau_{\delta})r_{\rm{LT},\delta})}-1\Big{)}\Big{)}\mathrm{d}u$
$\leq\int_{1-\gamma/2}^{\infty}\exp\Big{(}-\tilde{q}((1+\tau_{\delta})r_{\rm{LT},\delta})(1-\varepsilon)\Big{(}\big{(}\frac{u}{1+\tau_{\delta}}\big{)}^{\beta-\varepsilon}-1\Big{)}\Big{)}\mathrm{d}u$
$\rightarrow 0$
as $r_{\rm{LT},\delta}\to\infty$. We conclude that for every $\gamma>0$, the choice $y=y^{*}(1-\gamma)$ satisfies (32) for sufficiently small $\delta$, which implies that $\limsup_{\delta\downarrow 0}v_{{\bf CVaR},\delta}/v_{\delta}\geq 1-\gamma$. Theorem 2 now follows by letting $\gamma\downarrow 0$. $\square$
@@ 7.3. Proof of Theorem 3
Observe that $g(y)$ is strictly increasing by (9). The statement follows now immediately from Proposition 1. $\square$
@@ 7.4. Proof of Theorem 4
We recall that $r_{\delta}=\bar{F}^{-1}(\delta)$. We first derive the pointwise limit of ${\bf P}(\phi(y,L)>r_{\delta})/{\bf P}(|L|>r_{\delta})$ as $\delta\downarrow 0$, using (12). Let $\varepsilon\in(0,1/\max_{i}y^{T}A^{(i)}e)$, with $e$ a vector of ones. As we assumed that the entries of all $A^{(i)}$ are non-negative, $|L|\leq\varepsilon r_{\delta}$ implies that $\phi(y,L)\leq r_{\delta}$. Consequently,
$\frac{{\bf P}(\phi(y,L)>r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ $=\frac{{\bf P}(\phi(y,L)>r_{\delta};|L|>\varepsilon r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ (36)
$={\bf P}\big{(}\phi(y,L)>r_{\delta}\,\big{|}\,|L|>\varepsilon r_{\delta}\big{)}\frac{{\bf P}(|L|>\varepsilon r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ (37)
@@
Due to univariate regular variation, the ratio converges to $\varepsilon^{-\alpha}$. Due to the multivariate regular variation property (12), the conditional (on $|L|>\varepsilon r_{\delta}$) weak limit of $(L/|L|,|L|/r_{\delta})$ is $(\Theta,\varepsilon T)$, with $T$ a random variable with density $\alpha u^{-1-\alpha}$ independent of $\Theta$. Consequently,
$\frac{{\bf P}(\phi(y,L)>r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ $\rightarrow\varepsilon^{-\alpha}{\bf P}(\phi(y,\Theta)\varepsilon T>1)$ (39)
$=\varepsilon^{-\alpha}\int_{1}^{\infty}\alpha u^{-1-\alpha}{\bf P}(\phi(y,\Theta)\varepsilon>1/u)\mathrm{d}u$ (40)
@@
$=\int_{0}^{\infty}\alpha v^{\alpha-1}{\bf P}(\phi(y,\Theta)>v)\mathrm{d}v$ (42)
$={\bf E}[\phi(y,\Theta)^{\alpha}].$ (43)
In the derivation above, the second equality follows from the change of variable $v=1/\varepsilon u$, the third equality comes from the fact that $\phi(y,\Theta)\leq\max_{i}y^{T}A^{(i)}e\leq 1/\varepsilon$, and the last equality from a representation for the $\alpha$th moment of a non-negative random variable.
To finish the proof, we define $k=1/\delta$, $g_{k}(y)={\bf P}(\phi(y,L)>r_{\delta})/{\bf P}(|L|>r_{\delta})$ and $g(y)={\bf E}[\phi(y,\Theta)^{\alpha}]$, and apply Proposition 1. $\square$

### v2

@@ 8. Proofs
Let $g_{\ell},g:[0,\infty)^{m}\mapsto\mathbb{R}$ be functions, such that $y\mapsto g_{\ell}(y)$ is non-decreasing for each $\ell$, and $\eta\mapsto g(\eta y)$ is strictly increasing for each $y$ as a function of the scalar $\eta\in[0,\infty)$. Let $c\in[0,\infty)^{m}$, and define $v_{\ell}:=\sup\{c^{T}y:g_{\ell}(y)\leq 1,y\geq 0\}$ and $v_{\ast}:=\sup\{c^{T}y:g(y)\leq 1,y\geq 0\}$. Then, $v_{\ell}\to v_{\ast}$.
Proof.
Assume first that $v_{\ast}=\infty$. For each fixed $a>0$, there exists $y^{(a)}\geq 0$ such that $c^{T}y^{(a)}\geq a$ and $g(y^{(a)})\leq 1$. Fix $\eta\in(0,1)$. By the pointwise convergence $g_{\ell}(\eta y^{(a)})\to g(\eta y^{(a)})<1$, where the strict inequality follows from the fact that $g(\eta y^{(a)})$ is strictly increasing in $\eta$. In other words, $\eta y^{(a)}$ is feasible for $g_{\ell}$ for all sufficiently large $\ell$. Moreover, for such $\ell,v_{\ell}\geq c^{T}\eta y^{(a)}=\eta c^{T}y^{(a)}\geq\eta a$. Since $a$ is arbitrary, we obtain that $\liminf_{\ell\to\infty}v_{\ell}=\infty=v_{\ast}$. Assume now that $v_{\ast}<\infty$. Analogously to above, $v_{\ell}\geq\eta v_{\ast}$ for each $\eta\in(0,1)$ and all $\ell$ sufficiently large depending on $\eta$. As a result, $\liminf_{\ell\to\infty}v_{\ell}\geq v_{\ast}$. We turn to the upper bound.
Let us partition $c=(\bar{c},0)$, where the part $\bar{c}$ has strictly positive elements. We then note that $v_{*}=\sup\{\bar{c}^{T}\bar{y}:g(\bar{y},0)\leq 1,\bar{y}\geq 0\}$. Because $g_{\ell}$ is non-decreasing we have that $v_{\ell}=\sup\{\bar{c}^{T}\bar{y}:g_{\ell}(\bar{y},0)\leq 1,\bar{y}\geq 0\}.$ So, we may assume without loss of generality that $c$ has strictly positive elements.
Assume for contradiction that there exists $\varepsilon>0$ such that $(v_{\ell})_{\ell\geq 1}$ contains a subsequence such that $v_{\ell_{i}}\geq v_{\ast}+3\varepsilon$ for all $i$. Since $v_{\ell_{i}}=\sup\{c^{T}y:g_{\ell_{i}}(y)\leq 1,y\geq 0\}$, there must exist $y_{\ell_{i}}\in[0,\infty)^{m}$ such that $g_{\ell_{i}}(y_{\ell_{i}})\leq 1$ and $c^{T}y_{\ell_{i}}\geq v_{\ast}+2\varepsilon$. For each $i$, there exists $\eta_{i}\leq 1$ and $y^{\prime}_{\ell_{i}}\in[0,\infty)^{m}$ such that $y_{\ell_{i}}^{\prime}=\eta_{i}y_{\ell_{i}},g(y^{\prime}_{\ell_{i}})=g(\eta_{i}y_{\ell_{i}})\leq g(y_{\ell_{i}})\leq 1$ and $c^{T}y^{\prime}_{\ell_{i}}=v_{\ast}+2\varepsilon$.
Let $A:=\{y\in[0,\infty)^{n}:c^{T}y=v_{\ast}+2\varepsilon\}$. As all elements of $c$ are strictly positive, $A$ is a compact set. Define $f(y):=(\lfloor y_{1}\cdot|c|/\varepsilon\rfloor,\ldots,\lfloor y_{n}\cdot|c|/\varepsilon\rfloor)\cdot\varepsilon/|c|$, i.e., each coordinate is rounded down to the nearest multiple of $\varepsilon/|c|$. Define $\hat{A}:=\{f(y):y\in A\}\ni f(y^{\prime}_{\ell_{i}})=:\hat{y}_{\ell_{i}}$. Since $A$ is compact, $\hat{A}$ is a finite set, and thus we can extract a further subsequence such that $\hat{y}_{\ell_{i_{j}}}$ is identical to some $\hat{y}\in\hat{A}$ along the subsequence. By the rounding procedure, $c^{T}\hat{y}\geq v_{\ast}+\varepsilon$, and since $g_{\ell_{i_{j}}}$ is non-decreasing, $g_{\ell_{i_{j}}}(\hat{y})\leq 1$. By the pointwise convergence, $g(\hat{y})=\lim_{j\to\infty}g_{\ell_{i_{j}}}(\hat{y})\leq 1$, and thus $\hat{y}$ is a feasible solution for $g$. This contradicts that $v_{\ast}$ is the supremum of $\{c^{T}y:g(y)\leq 1,y\geq 0\}$. As a result, $\limsup_{\ell\to\infty}v_{\ell}\leq v_{\ast}+3\varepsilon$ for any $\varepsilon>0$. Thus, $\limsup_{\ell\to\infty}v_{\ell}\leq v_{\ast}$. $\square$
@@ 8.1. Proof of Theorem 1
We now turn to the proof of Theorem 1. We aim to apply Proposition 1 with $\ell=1/\delta$, $g_{\ell}(y)=(\log\delta)/\log{\bf P}(\phi(y,L)>q^{-1}(\log 1/\delta))$, and $g(y)=1/J(y)$. The pointwise convergence $g_{\ell}(y)\rightarrow g(y)$ follows from (10). The fact that $\eta\mapsto g(\eta y)$ is strictly increasing for all $y$ follows from the homogeneity property of $J$ implied by (9). Moreover, $\eta\mapsto g_{\ell}(\eta y)$ is non-decreasing because $\eta\mapsto\phi(\eta y,\ell)$ is non-decreasing for each $y$ and $\ell$. Thus, by Proposition 1, when $\delta$ is sufficiently small,
$q^{-1}(\log 1/\delta)\cdot v_{\delta}$ $=q^{-1}(\log 1/\delta)\cdot\sup\Big\{c^{T}x:{\bf P}\big(\phi(x,L)>1\big)\leq\delta,x\in X\Big\}$
$=\sup\Big\{c^{T}y:\delta\geq{\bf P}\big(\phi(y,L)>q^{-1}(\log 1/\delta)\big),y\geq 0\Big\}$
$=\sup\big\{c^{T}y:g_{\ell}(y)\leq 1,y\geq 0\big\}$
$\to\sup\big\{c^{T}y:g(y)\leq 1,y\geq 0\big\}=\sup\big\{c^{T}y:J(y)\geq 1,y\geq 0\big\}=c^{T}y^{\ast}.$
In the second equality above, we used that the constraint $x\in X$ is no longer relevant in ($\rm{LT}$) as $r\rightarrow\infty$, since we assumed $0,h]^{m}\subseteq X$ for some $h>0$, and $x=y/r$. $\square$
@@ 8.2. Proof of Theorem 2
Since $v_{{\bf CVaR},\delta}\leq v_{\delta}$, it suffices to prove an asymptotic lower bound. For this, we utilize Theorem 1. Let $y^{*}$ be a solution of ($\rm{LT}$) and for fixed $\gamma>0$, define $y:=(1-2\gamma)y^{*}$. Moreover, let $r=r_{\rm{LT},\delta}:=q^{-1}(\log 1/\delta)$. We will prove that $x=y/r_{\rm{LT},\delta}$ will be feasible for ($\rm{CVaR}_{\delta}$), for $\delta$ sufficiently small. We can write (5) with $x=y/r_{\rm{LT},\delta}$ as
${\bf CVaR}(\phi(y/r_{\rm{LT},\delta},L))=\inf_{\tau\in\mathbb{R}}\left[\tau+\frac{1}{\delta}\int_{\tau}^{\infty}{\bf P}(\phi(y,L)>ru)du\right]\leq 1-\gamma+\frac{1}{\delta}\int_{1-\gamma}^{\infty}{\bf P}(\phi(y,L)>ru)du.$ (28)
Using (10), for every $\varepsilon>0$ there exists an $r^{(1)}_{\varepsilon}$ such that
${\bf P}(\phi(y,L)>ru)=e^{-q(ru)J(y)(1+o(1))}\leq e^{-q(ru)J(y)(1-\varepsilon)},\qquad ru\geq r^{*}_{\varepsilon}.$ (29)
In addition, due to the Potter bounds (see e.g. Bingham et al. (1987), Theorem 1.5.6), there exists a constant $r^{(2)}_{\varepsilon}$ such that, for $u\geq 1/2$, and $r\geq r^{(2)}_{\varepsilon}$,
$q(ru)/q(r)\geq(1-\varepsilon)u^{\beta-\varepsilon}.$ (30)
Consequently, for $r$ sufficiently large and $u>1-\gamma$, using also that $J(y)=(1-2\gamma)^{-\beta}J(y^{*})=(1-2\gamma)^{-\beta}$,
${\bf P}(\phi(y,L)>ru)\leq e^{-q(r)u^{\beta-\varepsilon}(1-\varepsilon)^{2}/(1-2\gamma)^{\beta}}.$ (31)
Observe $\delta=\exp(-q(r))$, so that we get from (31) and (28)
${\bf CVaR}(\phi(y/r_{\rm{LT},\delta},L))\leq 1-\gamma+\int_{1-\gamma}^{\infty}e^{q(r)}e^{-q(r)u^{\beta-\varepsilon}(1-\varepsilon)^{2}/(1-2\gamma)^{\beta}}du$ (32)
Now, given $\gamma$, choose $\varepsilon$ small enough such that $(1-\gamma)^{\beta-\varepsilon}(1-\epsilon)^{2}/(1-2\gamma)^{\beta}>1$. In that case, the RHS of the last display converges to $1-\gamma$ as $r\to\infty$. We conclude that $x=(1-2\gamma)y^{*}/r$ is feasible for $\delta$ sufficiently small. Since this is true for each $\gamma>0$, the proof is complete. $\square$
@@ 8.3. Proof of Theorem 3
The functions $g_{k}(y)$ are non-decreasing for all $k$. We claim that $\eta\mapsto g(\eta y)$ is strictly increasing, which is implied if $I(y^{T}A^{(i)})$ is strictly decreasing for all $i=1,\ldots,d$. The latter is implied by the equality in the definition of $I$ in (9). The statement for arbitrary $k$ follows now immediately from Proposition 2.
Next, assume $k=k(\delta,\varepsilon)$ defined in (6). Then, as $\delta\downarrow 0$, for any $\varepsilon=\varepsilon(\delta)$ satisfying $\delta\log(1/\varepsilon)\to 0$,
$\log k(\delta,\varepsilon)$ $=\log\big((2/\delta)\log(1/\varepsilon)+2m+(2m/\delta)\log(2/\delta)\big)$
$=\log(1/\delta)+\log\big(2\log(1/\varepsilon)+2\delta/m+2m\log(2/\delta)\big)\sim\log(1/\delta).$ (35)
In view of Theorem 1 and the first part of Theorem 3, it suffices to show that, as $\delta\downarrow 0$,
$q^{-1}(\log 1/\delta)\sim q^{-1}(\log k(\delta,\varepsilon)),$ (36)
where $q(r)=h(r)r^{\beta}$, $\beta>1$, defined below (7). By Theorem 1.5.12 in Bingham et al. (1987), $q^{-1}$ is regularly varying with index $1/\beta$. The desired asymptotics in (36) follow now from (35). $\square$
@@ 8.4. Proof of Theorem 4
We recall that $r_{\delta}=\bar{F}^{-1}(\delta)$. We first derive the pointwise limit of ${\bf P}(\phi(y,L)>r_{\delta})/{\bf P}(|L|>r_{\delta})$ as $\delta\downarrow 0$, using (12). Let $\varepsilon\in(0,1/\max_{i}y^{T}A^{(i)}\mathrm{e})$, with $\mathrm{e}$ a vector of ones. As we assumed that the entries of all $A^{(i)}$ are non-negative, $|L|\leq\varepsilon r_{\delta}$ implies that $\phi(y,L)\leq r_{\delta}$. Consequently,
$\frac{{\bf P}(\phi(y,L)>r_{\delta})}{{\bf P}(|L|>r_{\delta})}=\frac{{\bf P}(\phi(y,L)>r_{\delta};|L|>\varepsilon r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ $={\bf P}\big(\phi(y,L)>r_{\delta}\,\big|\,|L|>\varepsilon r_{\delta}\big)\frac{{\bf P}(|L|>\varepsilon r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ (37)
@@
Due to univariate regular variation, the ratio converges to $\varepsilon^{-\alpha}$. Due to the multivariate regular variation property (12), the conditional (on $|L|>\varepsilon r_{\delta}$) distributional limit of $(L/|L|,|L|/r_{\delta})$ is $(\Theta,\varepsilon T)$, with $T$ a random variable with density $\alpha u^{-1-\alpha}$ independent of $\Theta$, see below (12) and Basrak et al. (2002). Consequently,
$\frac{{\bf P}(\phi(y,L)>r_{\delta})}{{\bf P}(|L|>r_{\delta})}\rightarrow\varepsilon^{-\alpha}{\bf P}(\phi(y,\Theta)\varepsilon T>1)$ $=\varepsilon^{-\alpha}\int_{1}^{\infty}\alpha u^{-1-\alpha}{\bf P}(\phi(y,\Theta)\varepsilon>1/u)\mathrm{d}u$ (39)
@@
$=\int_{0}^{\infty}\alpha v^{\alpha-1}{\bf P}(\phi(y,\Theta)>v)\mathrm{d}v={\bf E}[\phi(y,\Theta)^{\alpha}].$ (41)
[…]
To finish the proof, we define $k=1/\delta$, $g_{k}(y)={\bf P}(\phi(y,L)>r_{\delta})/{\bf P}(|L|>r_{\delta})$ and $g(y)={\bf E}[\phi(y,\Theta)^{\alpha}]$. By monotonicity of $y\mapsto\phi(y,L)$, these functions are all increasing. By Proposition 2, when $\delta$ is small,
$r_{\delta}v_{\delta}=r_{\delta}\sup\Big\{c^{T}x:{\bf P}\big(\phi(x,L)>1\big)\leq\delta,x\in X\Big\}$ $=\sup\Big\{c^{T}y:{\bf P}\big(\phi(y,L)>r_{\delta}\big)/{\bf P}(|L|>r_{\delta})\leq 1,y\geq 0\Big\}$
$\to\sup\Big\{c^{T}y:{\bf E}\big[\phi(y,\Theta)^{\alpha}\big]\leq 1,y\geq 0\Big\}=c^{T}y^{\ast}.$
We again used that the constraint $x\in X$ is replaced by $y\geq 0$, as we assumed $0,h]^{m}\subseteq X$ for some $h>0$, and $x=y/r$. $\square$